Consider the following.
θ = 30°, d = 20 m
(a) Find the length of the arc that subtends the given central angle θ on a circle of diameter d.
(b) Find the area A of the sector determined by θ.

Answer by Boreal(15235)

C=pi*d, and for a 30 degree arc, it would be (30/360)*pi*d=5.24 m, length of arc.
area is 1/12 pi*r^2 or 1/12 pi*100 cm^2=26.18 cm^3.